### v1

@@ 9.2 Appendix: Proofs
Now, by Assumption 18, the limit $\mathbf{V}^{\infty}:=\lim_{G\to\infty}\mathrm{E}[\mathbf{V}]$ exists. Besides, paralleling the existence argument in De Chaisemartin and D’Haultefoeuille (2020), one can show that $f$ is differentiable in a neighborhood of $\mathbf{V}^{\infty}$. Thus, the uniform Delta method (see, e.g., Van der Vaart, 1998, Theorem 3.8) yields:
@@
$\quad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
@@
$\qquad\qquad+\left(1-g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)\cdot\left(\left(\mathrm{E}\left[Q_{1,0,t}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t}^{*}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{*}\right]\right)\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\left.\left.+\left(\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{1,0,t}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{**}\right]\right)\right)\right)$
@@
where the the third equality follows from Assumption 7. By similar arguments:
@@
[…]
@@
$\quad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
@@
$\quad\ \mathrm{E}\left[P^{S}\right]=\mathrm{E}\left[\frac{N_{S^{*}}}{G}\right]=\mathrm{E}\left[\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S^{*}_{i,g,t}=1\}\right]$
@@
$\quad+\mathrm{E}\left[\frac{1}{G}\sum_{g=1}^{G}\mathbb{I}\{D_{g,T}=1,D_{g,T-1}=0\}\sum_{i=1}^{N_{g,T}}\mathbb{I}\{S^{*}_{i,g,T}=1\}\right]$
@@
$\quad\ \mathrm{E}\left[T^{S}\right]=\mathrm{E}\left[\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right]$
$=\frac{1}{G}\sum_{t\geq 2}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\left\{D_{g,t}=1,\ D_{g,t-1}=0\right\}\left(\mathrm{E}\left[\sum_{i=1}^{N_{g,t-1}}\mathbb{I}\left\{S_{i,g,t-1}^{\ast}=1\right\}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{D}_{g}\right]\right.\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \left.\left.+\mathrm{E}\left[\sum_{i=1}^{N_{g,t}}\mathbb{I}\left\{S_{i,g,t}^{\ast}=1\right\}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\mid\mathbf{D}_{g}\right]\right)\right]$
$=\frac{1}{G}\sum_{t\geq 2}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\left\{D_{g,t}=1,\ D_{g,t-1}=0\right\}\left(\mathrm{E}\left[\sum_{i=1}^{N_{g,t-1}}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid S_{i,g,t-1}^{\ast}=1,\mathbf{D}_{g}\right]\times\Pr\left(S_{i,g,t-1}^{\ast}=1\mid\mathbf{D}_{g}\right)\right.\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \left.\left.+\mathrm{E}\left[\sum_{i=1}^{N_{g,t}}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\mid S_{i,g,t}^{\ast}=1,\mathbf{D}_{g}\right]\times\Pr\left(S_{i,g,t}^{\ast}=1\mid\mathbf{D}_{g}\right)\right)\right]$
@@
where $\boldsymbol{\tau}_{1,0}$ was defined in (19). By Assumption 19, we have that $\mathrm{E}\left[\mathbf{U}(t,l)\right]$ converges pointwise for every $t$ and $l$ to $\mathbf{U}^{\infty}(t,l)$. Thus, also $\mathrm{E}\left[\mathbf{U}\right]$ converges to $\mathbf{U}^{\infty}$. Moreover, as $\boldsymbol{f}$ is differentiable in a neighborhood of $\mathbf{U}^{\infty}$, it follows again by the uniform Delta method (e.g., Theorem 3.8 Van der Vaart, 1998) that:
@@
and $\mathbf{0}_{4}$ denotes a 1×41$\times$ 41 × 4 vector of zeros. We thus have that:
@@
[…]
@@
Next, we show the first order asymptotic validity of the bootstrap moment equalities under the null hypothesis. Specifically, we want to show that under $H_{0}^{MC}$, for any $\nu>0$:
$\Pr\left(\sup_{\mathbf{x}\in\mathbb{R}^{\mathcal{C}_{S}}}\left|\operatorname*{\text{Pr}^{b}}\left(\sqrt{G}(\widehat{\boldsymbol{\tau}}_{1,0}^{b}-\widehat{\boldsymbol{\tau}}_{1,0})\leq\mathbf{x}\right)-\Pr\left(\sqrt{G}\widehat{\boldsymbol{\tau}}_{1,0}\leq\mathbf{x}\right)\right|>\nu\right)\rightarrow 0,$
@@
$\Pr\left(\sup_{x\in\mathbb{R}}\left|\operatorname*{\text{Pr}^{b}}\left(\sqrt{G}(\widehat{\tau}_{1,0}^{b}(t,l)-\widehat{\tau}_{1,0}(t,l))\leq x\right)-\Pr\left(\sqrt{G}\widehat{\tau}_{1,0}(t,l)\leq x\right)\right|>\nu\right)\rightarrow 0,$ (29)
@@
$=$ $\frac{1}{G}\sum_{g\in G}\left\{N_{g,t}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\left\{\left(Y_{g,t-1}-Y_{g,t-l}\right)\right.\right.$
@@
It remains to analyze the covariance of different elements of the vector $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}^{b}-\widehat{\boldsymbol{\tau}}\right)$. For simplicity, we only examine the terms involving (1,0)(1,0)( 1 , 0 ):
@@
×$\left.\times\frac{1}{P_{1,0}(j,k)\sqrt{G}}\sum_{g\in G}\left(\mathbb{I}\{D^{b}_{g,j}=1,D^{b}_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}^{b}-Y_{g,j-k}^{b}\right)-m(j,k)\right)\right]$
@@
×$\left.\times\left(\mathbb{I}\{D^{b}_{g,j}=1,D^{b}_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}^{b}-Y_{g,j-k}^{b}\right)-m(j,k)\right)\right]$
@@
×$\left.\times\left(\mathbb{I}\{D^{b}_{h,j}=1,D^{b}_{h,j-1}=0\}N_{h,j}\left(Y_{h,j-1}^{b}-Y_{h,j-k}^{b}\right)-m(j,k)\right)\right]$
@@
×$\times\left(Y_{g,t-1}-Y_{g,t-l}\right)\left(Y_{g,j-1}-Y_{g,j-k}\right)$
@@
×$\times\left(\frac{1}{P_{1,0}(j,k)G}\sum_{g\in G}\mathbb{I}\{D_{g,j}=1,D_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}-Y_{g,j-k}\right)\right),$
where the last equality follows since the blocks are, conditional on the sample, independent. We thus have that the bootstrap covariance is equal to the sample covariance, which converges in probability to $\lim_{G\rightarrow\infty}\mathrm{Cov}\left[\sqrt{G}(\widehat{\tau}_{1,0}(t,l)-\tau_{1,0}(t,l)),\sqrt{G}(\widehat{\tau}_{1,0}(j,k)-\tau_{1,0}(j,k))\right]$. Uniform convergence now follows by an application of the Cramér-Wold device, see for instance White (2000). Proposition 7 (i) then follows from continuous mapping.

### v2

@@ 9.2 Appendix: Proofs
Now, by Assumption 18, the limit $\mathbf{V}^{\infty}:=\lim_{G\to\infty}\mathrm{E}[\mathbf{V}]$ exists. Besides, paralleling the existence argument in CH2020, one can show that $f$ is differentiable in a neighborhood of $\mathbf{V}^{\infty}$. Thus, the uniform Delta method (see, e.g., VanderVaart1998, Theorem 3.8) yields:
@@
$\qquad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
@@
$\hskip 18.49988pt\hskip 18.49988pt+\left(1-g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)\cdot\left(\left(\mathrm{E}\left[Q_{1,0,t}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t}^{*}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{*}\right]\right)\right$.
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\qquad\left.\left.+\left(\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{1,0,t}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{**}\right]\right)\right)\right)$
@@
where the third equality follows from Assumption 7. By similar arguments:
@@
$\begin{aligned} \mathrm{E}\left[Y_{g,t}-Y_{g,t-1}\mid\mathbf{D}\right]&=-\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[\left(1-D_{i,g,t}^{*}\right)\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]\\ &\quad+\frac{1}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[\left(1-D_{i,g,t-1}^{*}\right)\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\mid\mathbf{D}\right]\\ &\quad+\mathrm{E}\left[Y_{1,t}(1)-Y_{1,t-1}(1)\right],\end{aligned}$
@@
$\qquad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
@@
$\qquad\ \mathrm{E}\left[P^{S}\right]=\mathrm{E}\left[\frac{N_{S^{*}}}{G}\right]=\mathrm{E}\left[\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S^{*}_{i,g,t}=1\}\right]$
@@
$\qquad+\mathrm{E}\left[\frac{1}{G}\sum_{g=1}^{G}\mathbb{I}\{D_{g,T}=1,D_{g,T-1}=0\}\sum_{i=1}^{N_{g,T}}\mathbb{I}\{S^{*}_{i,g,T}=1\}\right]$
@@
$\qquad\ \mathrm{E}\left[T^{S}\right]=\mathrm{E}\left[\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right]$
[…]
$\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\ \left.\left.+\mathrm{E}\left[\sum_{i=1}^{N_{g,t}}\mathbb{I}\left\{S_{i,g,t}^{\ast}=1\right\}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\mid\mathbf{D}_{g}\right]\right)\right]$
[…]
$\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\ \left.\left.+\mathrm{E}\left[\sum_{i=1}^{N_{g,t}}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\mid S_{i,g,t}^{\ast}=1,\mathbf{D}_{g}\right]\times\Pr\left(S_{i,g,t}^{\ast}=1\mid\mathbf{D}_{g}\right)\right)\right]$
@@
where $\boldsymbol{\tau}_{1,0}$ was defined in (19). By Assumption 19, we have that $\mathrm{E}\left[\mathbf{U}(t,l)\right]$ converges pointwise for every $t$ and $l$ to $\mathbf{U}^{\infty}(t,l)$. Thus, also $\mathrm{E}\left[\mathbf{U}\right]$ converges to $\mathbf{U}^{\infty}$. Moreover, as $\boldsymbol{f}$ is differentiable in a neighborhood of $\mathbf{U}^{\infty}$, it follows again by the uniform Delta method (e.g., Theorem 3.8 VanderVaart1998) that:
@@
and $\mathbf{0}_{4}$ denotes a $1\times 4$ vector of zeros. We thus have that:
@@
where $\mathbf{\Sigma}_{\mathbf{U}}$ is positive definite and $\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)$ has full rank, so also $\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)\boldsymbol{\Sigma}_{\mathbf{U}}\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)^{\prime}$ is positive definite. In particular, using $\infty$” also in the expectation operator to denote limiting expressions, e.g. $\mathrm{E}_{\infty}\left[Q_{0,0}(t,l)\right]:=\lim_{G\rightarrow\infty}\mathrm{E}\left[Q_{0,0}(t,l)\right]$, for a specific element $(t,l)$ from $\mathcal{C}_{S}$ we have that:
@@
Next, we show the first order asymptotic validity of the bootstrap moment equalities under the null hypothesis. Specifically, we want to show that, for any $\nu>0$:
$\Pr\left(\sup_{\mathbf{x}\in\mathbb{R}^{\mathcal{C}_{S}}}\left|\operatorname*{\text{Pr}^{b}}\left(\sqrt{G}(\widehat{\boldsymbol{\tau}}_{1,0}^{b}-\widehat{\boldsymbol{\tau}}_{1,0})\leq\mathbf{x}\right)-\Pr\left(\sqrt{G}(\widehat{\boldsymbol{\tau}}_{1,0}-\boldsymbol{\tau}_{1,0})\leq\mathbf{x}\right)\right|>\nu\right)\rightarrow 0,$
@@
$\Pr\left(\sup_{x\in\mathbb{R}}\left|\operatorname*{\text{Pr}^{b}}\left(\sqrt{G}(\widehat{\tau}_{1,0}^{b}(t,l)-\widehat{\tau}_{1,0}(t,l))\leq x\right)-\Pr\left(\sqrt{G}(\widehat{\tau}_{1,0}(t,l)-\tau_{1,0}(t,l))\leq x\right)\right|>\nu\right)\rightarrow 0,$ (29)
@@
[…]
@@
[…]
@@
$\left.\times\frac{1}{P_{1,0}(j,k)\sqrt{G}}\sum_{g\in G}\left(\mathbb{I}\{D^{b}_{g,j}=1,D^{b}_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}^{b}-Y_{g,j-k}^{b}\right)-m(j,k)\right)\right]$
@@
$\left.\times\left(\mathbb{I}\{D^{b}_{g,j}=1,D^{b}_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}^{b}-Y_{g,j-k}^{b}\right)-m(j,k)\right)\right]$
@@
$\left.\times\left(\mathbb{I}\{D^{b}_{h,j}=1,D^{b}_{h,j-1}=0\}N_{h,j}\left(Y_{h,j-1}^{b}-Y_{h,j-k}^{b}\right)-m(j,k)\right)\right]$
@@
$\times\left(Y_{g,t-1}-Y_{g,t-l}\right)\left(Y_{g,j-1}-Y_{g,j-k}\right)$
@@
$\times\left(\frac{1}{P_{1,0}(j,k)G}\sum_{g\in G}\mathbb{I}\{D_{g,j}=1,D_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}-Y_{g,j-k}\right)\right),$
where the last equality follows since the blocks are, conditional on the sample, independent. We thus have that the bootstrap covariance is equal to the sample covariance, which converges in probability to $\lim_{G\rightarrow\infty}\mathrm{Cov}\left[\sqrt{G}(\widehat{\tau}_{1,0}(t,l)-\tau_{1,0}(t,l)),\sqrt{G}(\widehat{\tau}_{1,0}(j,k)-\tau_{1,0}(j,k))\right]$. Uniform convergence now follows by an application of the Cramér-Wold device, see for instance White2000. Proposition 7 (i) then follows from continuous mapping.